# Demystifying Misconceptions in Social Bots Research

### v3

@@ 3 Conceptual issues
We see at least three strong arguments against this thesis. First, despite the limitations of bot detectors, there are several glaring examples of bot studies that were able to bring to light demonstrably harmful campaigns. For example, detectors developed as part of some scientific efforts were later deployed on online platforms and used to remove large numbers of malicious accounts (Yang et al., 2014). Similarly, the results of some bot activity studies led platforms to remove accounts identified as malicious bots (Ferrara, 2022; Yang and Menczer, 2023). In several other cases, scientific findings about bot activity were later found to be largely consistent with independent platform removals of malicious accounts (Nizzoli et al., 2020; Tardelli et al., 2022), confirming the accuracy of the scientific findings. These cases are just some of the success stories of social bot research. Without the development of bot detection techniques, identifying clusters of anomalous accounts engaged in sophisticated manipulations would not have been possible. Therefore, even if no universal bot detector exists yet, and despite the many caveats to consider in bot detection, being able to detect some malicious bots puts us in a more advantageous position than being able to detect none.
@@
Second, the methodological rigor and practical usefulness of some social bots studies are evidenced by corrective actions taken by government agencies based on academic research. For instance, in November 2014, the U.S. Securities and Exchange Commission (SEC) issued an alert to raise awareness about stock market manipulations exacerbated by social bots, [^7] Updated Investor Alert: Social Media and Investing – Avoiding Fraud (Nov. 12, 2014) https://www.sec.gov/resources-for-investors/investor-alerts-bulletins/ia_socialmediafraud following the findings of earlier studies (Hwang et al., 2012). Subsequent research confirmed these suspicions in multiple occasions (Cresci et al., 2019a; Nizzoli et al., 2020), contributing to calls for greater government oversight of social bots (Gorwa and Guilbeault, 2020a; Yan et al., 2023) and discussions on regulating bot freedom of expression (Lamo and Calo, 2019). Both the 2018 and 2022 EU Code of Practice on Disinformation, as well as investigations by the U.S. Congress and the UK Parliament’s Digital, Culture, Media and Sport Committee have also been informed by research on social bots. Furthermore, the concern over bot infiltration was highlighted when Elon Musk initially decided to abandon his purchase of Twitter, citing misleading information about the platform’s financial health and bot prevalence (Varol, 2023)—a claim also supported by Peiter Zatko, Twitter’s former head of security, who exposed serious deficiencies in the company’s security. [^8] What the Twitter Whistleblower Disclosure Means for Elon Musk: https://time.com/6207992/twitter-bots-elon-musk/ Table 2 provides some examples of contributions in the study and mitigation of social bots that have withstood the test of time. These examples underscore the enduring impact and significance of rigorous social bots research in shaping policy and industry practices.
@@
Third, the benefits of social bot research extend beyond the detection of malicious bots. For example, research and experimentation on social bots led to the development of neutral bots used to assess the level of political polarization and bias on a platform (Chen et al., 2021); “news bots” used for journalistic purposes to curate, aggregate, and distribute content gathered from multiple sources (Lokot and Diakopoulos, 2016); and even bots used for content moderation (Bilewicz et al., 2021; Askari et al., 2024). In addition, social bot research has contributed to the early development of other neighboring fields. The early research on social bots, which dates back to 2010 (Cresci, 2020), provided an important foundation, allowing the field to draw upon several years of accrued experience with related phenomena, thereby offering crucial insights when widespread concerns regarding misinformation, state-backed trolls, and coordinated inauthentic behavior surged in 2016 and subsequent years. In other words, early results on detecting and characterizing social bots informed strategies for detecting and mitigating other related forms of online manipulation. In light of these considerations, research on social bots—imperfect as it is—seems far from useless. In a sense, research on social bots is no exception to a core framework for the advancement of science: The accumulation of knowledge as a result of research efforts, other than solving only the specific problems the research is addressed to, also fertilizes the scientific ecosystem in many ways (e.g., forming generations of scientists) and provides tools, data, methodology, and insights for scientific advancements in close (and sometimes not so close) research fields.
@@ 4 Challenges in the post-API era
In addition to the widespread biases and misconceptions detailed above, the field of social bot research confronts significant obstacles due to changes in social media platform policies. In the API era, most social bot studies have focused on Twitter/X due to its free API and the ease of large-scale data acquisition. This hyperfocus on Twitter/X has led to the neglect of other platforms, creating a bias in social bots research that largely ignores platforms beyond Twitter/X. However, this data policy ceased in 2023 when Twitter/X terminated free data access for researchers. [^9] https://www.theverge.com/2023/5/31/23739084/twitter-elon-musk-api-policy-chilling-academic-research Similarly, Reddit restricted free data access, and Meta announced the discontinuation of CrowdTangle, a crucial tool for researchers studying Facebook and Instagram. [^10] https://www.niemanlab.org/2024/03/a-window-into-facebook-closes-as-meta-sets-a-date-to-shut-down-crowdtangle
@@
The lack of access to fresh social media data severely hampers researchers’ ability to monitor bot activities and assess their influence in real time. It significantly impedes the collection of new bot samples necessary for studying their characteristics and training novel machine learning classifiers. Even if researchers can develop new classifiers, deploying them at scale is challenging, exposing social media users to potential manipulation. Bot operators, on the other hand, remain largely unaffected. Leveraging burner and virtual cellphones, they can manage bot accounts across platforms directly, bypassing API restrictions. The advent of AI-powered social bots, which evade current bot detection models (Yang and Menczer, 2023; Yang et al., 2024), further emphasizes the importance of data availability in support of social bot research.
@@
Despite these setbacks, Europe’s Digital Services Act (DSA) has offered some hope by mandating large social media platforms to grant researchers data access upon reasonable requests. [^11] https://digital-strategy.ec.europa.eu/en/policies/digital-services-act-package Platforms like TikTok, Meta, and Google have initiated new data access programs. However, these programs are significantly limited compared to previous access levels. Their opaque and stringent application review processes also leave their efficacy in question (Jaursch et al., 2024). Besides traditional social media, the rise of decentralized platforms such as Mastodon and BlueSky introduce new dynamics. Their openness facilitates data access for researchers but also exposes them to exploitation by malicious actors. The decentralized nature of these platforms complicates efforts to combat threats such as malicious social bots, presenting novel challenges to both users and the research community. Moreover, the limited number of participants, which has not yet reached a critical mass, and the difficulty in creating a reliable ground truth further contributes to hindering bot detection research on these new platforms. Overall, the new data accessibility landscape and the change in the underlying technologies, represent an opportunity and a call for further research in social bot detection.
@@ 5 A call for moral responsibility
Our analysis of the drawbacks and limitations of recent studies brought to light the need for responsibility when discussing results about social bots. The way in which new findings are presented in this and in neighboring fields can influence not only the next iterations of research, but also industry practices, policymaking, and public opinion. To this end, it is paramount to avoid repeating the mistakes and propagating the common misconceptions that haunt the social bot literature and that currently fuel ambiguities, generalized misunderstandings, and friction among scholars.
The everyday challenges that we face as researchers in the broad area of misinformation are an accurate reflection of those faced by our society. As authors, reviewers, and readers of new research in this field, we have the moral obligation of refraining from falling for the same biases, and exacerbating the same issues, that we frequently encounter in our analyses. Misinformation often surfaces as «an accurate fact set in misleading context» (Starbird, 2019). Perpetuating the methodological and conceptual issues addressed in this paper contributes to creating flawed or unreliable research. Cherry-picking data, references, claims, or results from a cited article are examples of misleading context in which misinformation thrives (West and Bergstrom, 2021). Similarly, while making hyped or sensationalist claims can help publish a paper and accrue citations, such claims also contribute to creating misleading contexts and unrealistic expectations, as well as to increasing scientific tribalism.
@@ 6 Acknowledgments
This work was partially supported by project SERICS (PE00000014) under the NRRP MUR program funded by the EU – NGEU; by the European Union – Horizon 2020 Program under the scheme “INFRAIA-01-2018-2019 – Integrating Activities for Advanced Communities”, Grant Agreement n. 871042, “SoBigData++: European Integrated Infrastructure for Social Mining and Big Data Analytics”; by project DESIRE 2.0 (DissEmination of ScIentific REsults) funded by IIT-CNR; by the Volkswagen Foundation as part of the Bots Building Bridges project, under the “Artificial Intelligence and the Society of the Future” initiative; by DARPA (awards W911NF-17-C-0094 and HR001121C0169); and by the Knight Foundation.

### v4

@@ 3 Conceptual issues
We see at least three strong arguments against this thesis. First, despite the limitations of bot detectors, there are several glaring examples of bot studies that were able to bring to light demonstrably harmful campaigns. For example, detectors developed as part of some scientific efforts were later deployed on online platforms and used to remove large numbers of malicious accounts (Yang et al., 2014). Similarly, the results of some studies led platforms to remove accounts identified as malicious bots (Ferrara, 2022; Yang and Menczer, 2023). In several other cases, scientific findings about bot activity were later found to be consistent with independent platform removals of malicious accounts (Nizzoli et al., 2020; Tardelli et al., 2022), confirming the accuracy of the scientific findings. These cases represent just some of the success stories of social bot research. Therefore, even if no universal bot detector exists, and despite the many caveats to consider in bot detection, being able to detect some malicious bots puts us in a more advantageous position than being able to detect none.
@@
Second, the methodological rigor and practical usefulness of some social bots studies are evidenced by corrective actions taken by government agencies based on academic research. For instance, in November 2014, the U.S. Securities and Exchange Commission (SEC) issued an alert to raise awareness about stock market manipulations exacerbated by social bots, [^6] Updated Investor Alert: Social Media and Investing – Avoiding Fraud (Nov. 12, 2014) https://www.sec.gov/resources-for-investors/investor-alerts-bulletins/ia_socialmediafraud following the findings of earlier studies (Hwang et al., 2012). Subsequent research confirmed these suspicions on multiple occasions (Cresci et al., 2019a; Nizzoli et al., 2020), contributing to calls for greater government oversight of social bots (Gorwa and Guilbeault, 2020; Yan et al., 2023) and discussions on regulating bot freedom of expression (Lamo and Calo, 2019). Both the 2018 and 2022 EU Code of Practice on Disinformation, as well as investigations by the U.S. Congress and the UK Parliament’s Digital, Culture, Media and Sport Committee, have also been informed by research on social bots. Furthermore, the concern over bot infiltration was highlighted when Elon Musk initially decided to abandon his purchase of Twitter, citing misleading information about the platform’s financial health and bot prevalence (Varol, 2023)—a claim also supported by Peiter Zatko, Twitter’s former head of security, who exposed serious deficiencies in the company’s security. [^7] What the Twitter Whistleblower Disclosure Means for Elon Musk: https://time.com/6207992/twitter-bots-elon-musk/
@@
Third, the benefits of social bot research extend beyond the detection of malicious bots. For example, research and experimentation on social bots led to the development of neutral bots used to assess the level of political polarization and bias on a platform (Chen et al., 2021); “news bots” used for journalistic purposes to curate, aggregate, and distribute content gathered from multiple sources (Lokot and Diakopoulos, 2016); and even bots used for content moderation (Bilewicz et al., 2021; Askari et al., 2024). In addition, social bot research has contributed to the early development of other neighboring fields. The early research on social bots, which dates back to 2010 (Cresci, 2020), provided an important foundation, allowing the field to draw upon several years of accrued experience when widespread concerns regarding misinformation, state-backed trolls, and coordinated inauthentic behavior surged in 2016 and subsequent years. In other words, early results on detecting and characterizing social bots informed strategies for detecting and mitigating other related forms of online manipulation. In light of these considerations, research on social bots—imperfect as it is—seems far from useless. In fact, social bots research perfectly exemplifies the process leading the general advancement of science: the accumulation of knowledge resulting from certain research efforts, other than solving local problems, also fertilizes the broader scientific ecosystem, providing data, methodologies, tools, and insights for further scientific advancements in close—and sometimes not so close—fields.
@@ 4 Challenges in the post-API era
In addition to the widespread biases and misconceptions detailed above, the field of social bot research confronts significant obstacles due to changes in social media platform policies. In the API era, most social bot studies have focused on Twitter/X due to its free API and the ease of large-scale data acquisition. This hyperfocus on X has led to the neglect of other platforms, creating a bias in social bots research that largely ignores other platforms. However, this data policy ceased in 2023 when X terminated free data access for researchers. [^8] https://www.theverge.com/2023/5/31/23739084/twitter-elon-musk-api-policy-chilling-academic-research Similarly, Reddit restricted free data access, and Meta announced the discontinuation of CrowdTangle, a crucial tool for researchers studying Facebook and Instagram. [^9] https://www.niemanlab.org/2024/03/a-window-into-facebook-closes-as-meta-sets-a-date-to-shut-down-crowdtangle
@@
The lack of access to fresh social media data severely hampers researchers’ ability to monitor bot activities and assess their influence. It significantly impedes the collection of new bot samples necessary for studying their characteristics and training novel machine learning classifiers. Even if researchers can develop new classifiers, deploying them at scale is challenging, exposing social media users to potential manipulation. Bot operators, on the other hand, remain largely unaffected. Leveraging burner and virtual cellphones, they can manage bot accounts across platforms directly, bypassing API restrictions. The advent of AI-powered social bots, which evade current bot detection models (Yang and Menczer, 2023; Yang et al., 2024), further emphasizes the importance of data availability in support of social bot research.
@@
Despite these setbacks, Europe’s Digital Services Act (DSA) has offered some hope by mandating large social media platforms to grant researchers data access upon reasonable requests. [^10] https://digital-strategy.ec.europa.eu/en/policies/digital-services-act-package Platforms like TikTok, Meta, and Google have initiated new data access programs. However, these programs are significantly limited compared to previous access levels. Their opaque and stringent application review processes also leave their efficacy in question (Jaursch et al., 2024). Besides traditional social media, the rise of decentralized platforms such as Mastodon and BlueSky introduces new dynamics. Their openness facilitates data access for researchers but also exposes them to exploitation by malicious actors. The decentralized nature of these platforms complicates efforts to combat threats such as malicious social bots, presenting novel challenges to both users and the research community. Moreover, the limited number of participants, which has not yet reached a critical mass, and the difficulty in creating a reliable ground truth, further contribute to hindering bot detection research on these platforms. Overall, the new data accessibility landscape and the change in the underlying technologies, represent an opportunity and a call for further research in social bot detection.
@@ 5 A call for moral responsibility
Our analysis of the drawbacks and limitations of recent studies brought to light the need for responsibility when discussing results about social bots. The way in which new findings are presented in this and in neighboring fields can influence not only the next iterations of research, but also industry practices, policymaking, and public opinion. To this end, it is paramount to avoid repeating the mistakes and propagating the common misconceptions that currently haunt the social bot literature and that fuel ambiguities, generalized misunderstandings, and friction among scholars.
The everyday challenges that we face as researchers in the broad area of misinformation are an accurate reflection of those faced by our society. As authors, reviewers, and readers of new research in this field, we have the moral obligation of refraining from falling for the same biases, and exacerbating the same issues, that we frequently encounter in our analyses. Misinformation often surfaces as «an accurate fact set in a misleading context» (Starbird, 2019). Perpetuating the methodological and conceptual issues addressed in this paper contributes to creating flawed or unreliable research. Cherry-picking data, references, claims, or results from a cited article are examples of misleading context in which misinformation thrives (West and Bergstrom, 2021). Similarly, while making hyped or sensationalist claims can help publish a paper and accrue citations, such claims also contribute to creating misleading contexts and unrealistic expectations, as well as to increasing scientific tribalism.
@@ 6 Acknowledgments
This work was partially supported by the European Union – Next Generation EU, Mission 4 Component 1, for project PIANO (CUP B53D23013290006) and for the ERC project DEDUCE under grant #101113826.; by project DESIRE 2.0 (DissEmination of ScIentific REsults) funded by IIT-CNR; by the Volkswagen Foundation as part of the Bots Building Bridges project, under the “Artificial Intelligence and the Society of the Future” initiative; by DARPA (awards W911NF-17-C-0094 and HR001121C0169); and by the Knight Foundation.